OpenStudy (anonymous):

Jen and Amber went walking. Jenn has a 20 sec head start. Jen walks at 4 ft per sec. Amber walks at 6 ft per second. How many seconds will Jenn have been walking when they walk the same distance.

OpenStudy (anonymous):

I know Jenn has an 80 foot head start. I dont know if the equation is y+80=6x and y=4x.

OpenStudy (anonymous):

20+4t=y 6t=y

OpenStudy (anonymous):

or 80+4t=y and 6t=y

OpenStudy (anonymous):

80+4t=6t 80=2t 40=t

OpenStudy (anonymous):

oh yay an answer--hang on

OpenStudy (anonymous):

I was watching a math teacher video on equation eliminations...thank you. on this one it is really the equation that was stumping me.

OpenStudy (anonymous):

but isnt the 80 a red herring...that is feet. distance =time x rate. Is it the 80 or the 20 that we are adding on...the answer is to be in seconds, not feet.